# Investigating the Representation of Backchannels and Fillers in Fine-tuned Language Models

### v1

@@ D.5 Confusion Matrices for the Top 15 Backchannels and Fillers in English and Japanese
yeah 3 2 4 4 2 2 4 2
uh-huh 2 2 5 3 2 2 4 2
well 2 2 3 4 2 2 2 2
right 2 2 6 6 2 2 2 2
oh 2 2 3 3 2 2 2 2
um 2 2 2 2 2 2 3 3
okay 2 2 2 2 2 2 2 2
@@
Average 2.13 2.00 3.00 2.87 2.00 2.00 2.67 2.20
Table 18: K values of optimal results on the k-means clustering of the top 15 selected Japanese backchannels/fillers when using the LLaMA-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
no_ft_one_ctx
ft_one_ctx
(next-token)
no_ft_full_ctx
ft_full_ctx
(next-token)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Table 19: K values of optimal results on the k-means clustering of the top 15 selected Japanese backchannels/fillers when using the Qwen-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx ft_no_ctx no_ft_one_ctx ft_one_ctx no_ft_full_ctx ft_full_ctx
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Table 20: K values of optimal results on the k-means clustering of the top 15 selected Japanese backchannels/fillers when using the GPT-2 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
ft_no_ctx
(turn-taking)
no_ft_one_ctx
ft_one_ctx
(next-token)
ft_one_ctx
(turn-taking)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Table 21: K values of optimal results on the k-means clustering of the top 15 selected Japanese backchannels/fillers when using the BERT model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx
ft_no_ctx
no_ft_one_ctx
ft_one_ctx
ori 100 ori 100 ori 100 ori 100
うん (un) 3 3 3 3 2 3 2 2
あ (a) 2 4 4 3 3 3 4 3
はい (hai) 2 2 3 3 2 2 5 6
え (e) 4 2 3 4 5 5 4 4
そう (sou) 2 2 2 2 2 2 2 3
ま (ma) 4 4 3 2 4 3 3 3
なんか (nanka) 2 2 2 2 2 2 3 3
あの (ano) 2 3 3 3 2 2 3 2
@@
いや (iya) 2 2 3 3 3 2 3 3
へー (he-) 2 2 2 2 2 2 2 2
そうか (souka) 2 2 3 2 2 2 2 2
Average 2.40 2.47 2.87 2.87 2.47 2.40 3.00 3.00

### v2

@@ D.7 Additional Confusion Matrices
yeah 2 3 2 2 2 2 3 3
uh-huh 3 3 3 2 2 2 2 2
well 2 2 2 3 2 2 2 2
right 2 2 2 2 2 2 2 2
oh 4 3 2 2 2 2 3 5
um 2 2 2 2 2 2 2 2
okay 9 9 2 4 2 2 2 2
@@
Average 3.00 2.80 2.20 2.33 2.06 2.20 2.20 2.27
Table 19: $k$-values of optimal results on the $k$-means clustering of the top 15 selected Japanese backchannels/fillers when using the LLaMA-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Table 20: $k$-values of optimal results on the $k$-means clustering of the top 15 selected Japanese backchannels/fillers when using the Qwen-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Table 21: $k$-values of optimal results on the $k$-means clustering of the top 15 selected Japanese backchannels/fillers when using the GPT-2 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) ft, no_ctx (TTP) no_ft, one_ctx ft, one_ctx (NTP) ft, one_ctx (TTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Table 22: $k$-values of optimal results on the $k$-means clustering of the top 15 selected Japanese backchannels/fillers when using the BERT model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (MASK) no_ft, one_ctx ft, one_ctx (MASK)
orig. 100 orig. 100 orig. 100 orig. 100
うん (un) 2 2 2 3 4 2 3 3
あ (a) 3 4 4 2 3 4 4 2
はい (hai) 2 2 2 4 2 2 2 2
え (e) 4 4 2 3 5 6 2 3
そう (sou) 2 2 2 2 2 4 2 3
ま (ma) 4 4 2 2 4 4 2 3
なんか (nanka) 2 2 2 2 2 2 2 2
あの (ano) 2 2 2 2 3 3 2 2
@@
いや (iya) 4 2 2 2 3 2 2 2
へー (he-) 2 2 2 2 2 2 2 4
そうか (souka) 3 3 2 2 2 2 2 2
Average 2.60 2.53 2.20 2.47 2.66 2.73 2.46 2.73
(a) no_ft, one-context
(b) ft, one-context (NTP)
Figure 21: Distance matrices for the top 15 English backchannels/fillers in the LLaMA-3 model (a) before and (b) after fine-tuning.
(a) no_ft, one-context
(b) ft, one-context (NTP)
Figure 22: Distance matrices for the top 15 English backchannels/fillers in the Qwen-3 model (a) before and (b) after fine-tuning.
(a) no_ft, one-context
(b) ft, one-context (NTP)
Figure 23: Distance matrices for the top 15 English backchannels/fillers in the GPT-2 model (a) before and (b) after fine-tuning.
(a) no_ft, one-context
(b) ft, one-context (MASK)
Figure 24: Distance matrices for the top 15 English backchannels/fillers in the BERT model (a) before and (b) after fine-tuning.
(a) no_ft, one-context
(b) ft, one-context (NTP)
Figure 25: Distance matrices for the top 15 Japanese backchannels/fillers in the LLaMA-3 model (a) before and (b) after fine-tuning.
(a) no_ft, one-context
(b) ft, one-context (NTP)
Figure 26: Distance matrices for the top 15 Japanese backchannels/fillers in the GPT-2 model (a) before and (b) after fine-tuning.
(a) no_ft, one-context
(b) ft, one-context (MASK)
Figure 27: Distance matrices for the top 15 Japanese backchannels/fillers in the BERT model (a) before and (b) after fine-tuning.